OpenStudy (anonymous):

Following the recommended steps for solving equations, what should be your first step in solving the given equation? 2x – 3 – 4x = –6 + 2x A. add 3 to both sides B. subtract 4x from 2x C. subtract 2x from both sides D. add 4x to 2x

OpenStudy (anonymous):

2x-3-4x+3=-6+3+2x 2x-4x=-2x -3+3=0 -6+3=-3 therefore equation is -2x=2x-3 adding 2x on both sides 2x-2x=2x+2x-3 4x-3=0 4x=3 x=3/4

OpenStudy (anonymous):

add 3 to both sides is what I would do

OpenStudy (anonymous):

that is what i did

OpenStudy (anonymous):

okay now im lost aren't you supposed to combine the like terms?

OpenStudy (anonymous):

if you follow the steps you have to but I take the easy route and do whatever looks easiest first

OpenStudy (anonymous):

oh alright because having a lot of variables and numbers,having to combine them confuses me A LOT

OpenStudy (anonymous):

I would combine like terms first, it does make it look better to solve if you're a newbie at equations.

OpenStudy (anonymous):

alright thank you so much for helping me

OpenStudy (anonymous):

:) no problem